A restaurant chef mixes 20 lb. of snow peas costing $1.99 a pound with 14 lb. of petite onions costing $1.19 a pound to make a vegetable medley for the evening meal. Find the cost per pound of the mixture. (You must show alg. to receive full credit)
Found 2 solutions by checkley77, 24HoursTutor.com:
Answer by checkley77(12844)

20*1.99+14*1.19=34x
39.8+16.66=34x
56.46=34x
x=56.46/34
x=$1.66 is the price of the mix.

Answer by 24HoursTutor.com(40)

Rate of 1 pound of snow peas = $1.99
Cost of 20 lbs. of snow peas = 1.99 X 20 = 39.8 $
Rate of 1 pound of petite DISABLED_event_onions= $1.19
Cost of 14 lbs. of petite DISABLED_event_onions= 1.19 X 14 = 16.66 $
Total weight of vegetable medley = weight of snow peas + weight of petite onions
= 20 + 14
= 36 Lbs.
Total cost of 36 lbs. of vegetable medley = cost of snow peas + cost of onions
= 39.80 $ + 16.66 $
= 56.46 $
Total cost of 1 pound of vegetable medley = 56.46 / 36
= 1.57 $
Ans.: The cost of per pound of the vegetable medley would be $1.57
The above question is solved by one of the experts from 24HoursTutor.com
We provide live 1:1 online homework help
1. 24X7 tutors available
2. Certified tutors
3. Only for $9.99/hr